OpenStudy (rane):

The purity os sodium hydrogen carbonate sample can be found bye experiment. In such a exercise, Mona added a carefully weighed NaHCO3 sample to a conical flask containing a concentrated hydrochloric solution. the data recorded is shown here: most of impure NaHCO3...... 4.262g Mass of flask + HCL solution ... 57.982g mass of flask + contents after the reaction was complete.........60.370g find the mass of NaHCO3 actually present in the impure measured sample.

OpenStudy (rane):

@chmvijay

OpenStudy (rane):

u guys might think im stupid but at this time - midnight- especially after doing maths for 5 hrs- i have nothing in my brain left. so guys please co-operate here

OpenStudy (chmvijay):

@RANE by the way who is mona

OpenStudy (rane):

i dont know its just what the question states, she is my class mate to be honest but the teacher choose to use her name zoo

OpenStudy (chmvijay):

balanced equation ?

OpenStudy (chmvijay):

mass of flask + contents after the reaction was complete.........60.370g is that including the Gas CO2 liberated in the reaction or what ?

OpenStudy (rane):

this is equation but its no balancing NaHCO3 + HCl = H20 + CO2 + NaCl

OpenStudy (rane):

wait its already balanced

OpenStudy (chmvijay):

60.370g is that including the Gas CO2 liberated in the reaction or what ?

OpenStudy (rane):

idk this is all the information given, which i have posted above

OpenStudy (chmvijay):

i consider that they have weighed CO2 also and will try to solve it OK

OpenStudy (chmvijay):

most of impure NaHCO3+Mass of flask + HCL solution = 57.982g +4.262g =62.244 gram but the total mass of the product is = 60.370g so for 62.224 gram we are getting = 60.370 g for 100 gram = 60.370 *100 /62.244 = claculate that is ur purity

OpenStudy (rane):

vijay i did completely different to what u did

OpenStudy (anonymous):

hmmm m fine ... thanks .. wt about u

OpenStudy (chmvijay):

what u did let me see that

OpenStudy (chmvijay):

iam fine musakan rane told me that ur her good friend

OpenStudy (anonymous):

yea she is my best frend

OpenStudy (rane):

n(NaHC03) = 0.045mols n(NaHCO3) -actual [resent = 1/1*0.045mols mass = 0.0045*9 = 4.0495g

OpenStudy (chmvijay):

good :)

OpenStudy (rane):

lolz muskan ab hasti na rehna pls

OpenStudy (anonymous):

hahha .. :D

OpenStudy (chmvijay):

according to ur answer its 95 % pure and from my answer its 97 % pure :) not much difference:) how you got 0.045mols

OpenStudy (chmvijay):

nice smile @Muskan

OpenStudy (anonymous):

tabhi nam muskan rakha hai .... :) q @RANE

OpenStudy (rane):

n = m/M = 4.262/91 = 0.045moles

OpenStudy (rane):

hahah very smart @Muskan lakin mairy liyai to tum khuch aur hi ho

OpenStudy (anonymous):

ahahha ... may tumhare liye tumhari DARLING hun u 9 :D

OpenStudy (rane):

hahahah u never know

OpenStudy (anonymous):

hahah .. chalo bata do kia hun tumhare liye may

OpenStudy (chmvijay):

i dont know really rane ur on other end of my calculation whcih i hat e a lot that is moles :P

OpenStudy (rane):

lol i never knew u hated something in chemistry SHOCKING !!!

OpenStudy (anonymous):

i also hate NaHCo3, H2SO4, CL2 etc.................................

OpenStudy (rane):

thats the ones i love but atm i hate everything just wana go to bed and sleep

OpenStudy (anonymous):

aur tumhe pata hai agy chemical eng karni hai may :_D

OpenStudy (rane):

inshaalah lolz

OpenStudy (anonymous):

hahaha tab tk sab se piayar ho jae ga

OpenStudy (rane):

ha! we'll se "sab sy" !!!!

OpenStudy (anonymous):

tum se to pehly din se hai ... baki sab means NaHCO3, CO2, H2O.....etc

OpenStudy (rane):

hahaha

OpenStudy (chmvijay):

n(NaHC03) = 0.045mols n(NaHCO3) -actual [resent = 1/1*0.045mols mass = 0.0045*9 = 4.0495g how u got this ?

OpenStudy (anonymous):

:)